OpenStudy (anonymous):

A certain store has a fax machine available for use by its customers. The store charges $2.30 to send the first page and $0.40 for each subsequent page. The total price, P, for the faxing x pages can be modeled by the formula . Determine the number of pages that can be faxed for $5.90. A) 10 pages B) 15 pages C) 3 pages D) 55 pages

OpenStudy (anonymous):

sweet, can you write the formula?

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

p=0.40(x-1)+2.30

OpenStudy (anonymous):

number four on the review

OpenStudy (anonymous):

what is p?

OpenStudy (anonymous):

i cant believe i forgot how to do this.

OpenStudy (anonymous):

plug in the numbers and u r good

OpenStudy (anonymous):

so 5.90=0.40(40-1)=2.30

OpenStudy (anonymous):

oops that +2.30

OpenStudy (anonymous):

yes whats next?

OpenStudy (anonymous):

wait where did the 40 come from?

OpenStudy (anonymous):

40 cents = x ?

OpenStudy (anonymous):

or no

OpenStudy (anonymous):

10 pages

OpenStudy (anonymous):

no u have the formula already P = 0.40*(x-1)+2.30 p is the total price 0.40 the 40 cents x is the number of papers -1 minus the 1st page 2.30 the first page price

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i plugged 10 in to x

OpenStudy (anonymous):

.40(10)= 4-.4+2.30=5.90 :D

OpenStudy (anonymous):

5.90 = 0.40*(x-1)+2.30 solve for x

OpenStudy (anonymous):

moving on thank you

OpenStudy (anonymous):

5.90 = 0.40*(x-1)+2.30 1st subtract a 2.30 from both sides

OpenStudy (anonymous):

3.6

OpenStudy (anonymous):

write the whole thing u kno i love to make u do extra work

OpenStudy (anonymous):

nope u just have to live with it or get your butt to the other question.